# Nearly-Linear Time Seeded Extractors with Short Seeds

### v1

@@ 3.1 Fast Generation of Small-Bias Sets
A set $S\subseteq\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}$ is $\varepsilon$-biased if the uniform distribution over its elements is indistinguishable from uniform by every linear test. Namely, if for every nonempty $T\subseteq[n]$ it holds that $\Pr_{s\sim S}[\bigoplus_{i\in T}s_{i}]\in[\frac{1-\varepsilon}{2},\frac{1+\varepsilon}{2}n]$. We say that a linear code $\mathcal{C}\subseteq\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}$ is $\varepsilon$-balanced if the Hamming weight of each nonzero codeword lies in $[\frac{1-\varepsilon}{2}n,\frac{1+\varepsilon}{2}]$. It is known that these two objects are essentially the same: $S$ is $\varepsilon$-biased if and only if the $|S|\times n$ matrix whose rows are the elements of $S$ is a generator matrix of an $\varepsilon$-balanced code.
@@
Lemma 3.1 (RWs amplify bias [Ta-17] [^10] The argument for $t=2$ was suggested already by Rozenman and Wigderson (see [Bog12]).).
@@
There exist constants $R\in(0,1)$ and $\varepsilon_{0}\in(0,1)$ such that there exists an explicit family of codes $\mathopen{}\mathclose{{}\left\{{\mathrm{Jus}_{n}}}\right\}$, each of which has block length $n$, rate $R$, and is $\varepsilon_{0}$-balanced. Moreover, given $x\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{k=Rn}$, $\mathrm{Jus}_{n}(x)$ can be computed in $\widetilde{O}(n)$.
@@
The parameters of the codes follow from the original construction (and specifically, the lemma holds, say, with $R=\frac{1}{8}$ and $\varepsilon_{0}=\frac{37}{40}$), so we will just show that the code is efficiently computable. Given a message $x$, we first encode it with a full Reed–Solomon code of constant relative distance over a field $\mathds{F}_{q}$ of characteristic $2$, where $q\log q=O(n)$. By Lemma 2.1, this can be done in time $\widetilde{O}(q)=\widetilde{O}(n)$. Then, we encode with polynomial evaluation $p_{x}(\alpha)$, for $\alpha\in\mathds{F}_{q}$, with the binary representation of $(p(\alpha),\alpha\cdot p(\alpha))$. This takes $\widetilde{O}(q)$ time as well. ∎
@@
There exist a constant $c>1$, and an explicit family of balanced codes, such that for every $\bar{n}\in\mathbb{N}$ and any $\varepsilon>0$, $\mathcal{C}\subseteq\mathds{F}_{2}^{\bar{n}}$ is $\varepsilon$-balanced of rate $R=\varepsilon^{c}$, and given $x\in\mathds{F}_{2}^{k=R\bar{n}}$, any $m$ bits of $\mathcal{C}(x)$ can be computed in time $\widetilde{O}(n)+O(m\log(1/\varepsilon)\log n\operatorname{loglog}n)$.
Moreover, for every $k\in\mathbb{N}$ and any $\varepsilon>0$ there exists an explicit $\varepsilon$-biased set over $\mathds{F}_{2}^{k}$ generated by a function $\mathsf{SmallBias}\colon[\bar{n}]\rightarrow\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{k}$ computable in time $(k+\log(1/\varepsilon))\cdot\widetilde{O}(\log k)$.
@@
Let $\mathcal{C}_{0}\colon\mathds{F}_{2}^{k=\Theta(n)}\rightarrow\mathds{F}_{2}^{n}$ be the $\varepsilon_{0}$-balanced code guaranteed to us by Lemma 3.2, and let $G=(V=[n],E)$ be the $D$-regular $\lambda$-spectral expander of Corollary 2.18, instantiated with $\lambda=\frac{1-\varepsilon_{0}}{4}$ (so $D=D(\varepsilon_{0})$). Letting $\mathcal{C}\colon\mathds{F}_{2}^{k}\rightarrow\mathds{F}_{2}^{\bar{n}}$ be the amplified code of Lemma 3.1 set with
$t=\frac{2\log(1/\varepsilon)}{\log\mathopen{}\mathclose{{}\left(\frac{1+\varepsilon_{0}}{2}}\right)}=O(\log(1/\varepsilon)),$
the lemma tells us that it is $(\varepsilon_{0}+2\lambda)^{t/2}\leq\varepsilon$ balanced. Given $x\in\mathds{F}_{2}^{n}$ and $i\in[\bar{n}]$, computing $\mathcal{C}(x)_{i}$ amounts to XORing $t$ coordinates of $\mathcal{C}_{0}(x)$ determined by $i=(v,i_{1},\ldots,i_{t})$, which indexes a random walk over $G$. Computing $\mathcal{C}_{0}(x)$ takes $\widetilde{O}(n)$ time, and computing a length-$t$ walk over $G$ takes $t\cdot O(\log(1/\lambda)\cdot\log n\cdot\operatorname{loglog}n)$ time. The corollary then follows, observing that $\bar{n}=n\cdot D^{t}=n\cdot\operatorname{poly}(1/\varepsilon)$, and that
$\widetilde{O}(n)+m\cdot t\cdot O(\log n\cdot\operatorname{loglog}n)=\widetilde{O}(n)+m\cdot\log(1/\varepsilon).$
@@
$t\cdot O(\log(1/\lambda)\cdot\log n\cdot\operatorname{loglog}n)+k\cdot\widetilde{O}(\log n)=(k+\log(1/\varepsilon))\cdot\widetilde{O}(\log k).\qed$
@@
Instead of using Justesen’s code from Lemma 3.2 as our inner code $\mathcal{C}_{0}$, we can instead use the linear-time encodable code of Spielman [Spi96]. While not stated as balanced codes, but rather as constant relative distance codes, one can verify that the distance can also be bounded by above. The construction is more involved than Justesen’s one. However, in the logarithmic-cost RAM model, in which basic register operations over $O(\log n)$ bit registers count as a single time step, Spielman’s code can be implemented in $O(n)$ time.
@@ 3.2 A Sampler from Bounded Independence
Recall that $X_{1},\ldots,X_{n}\sim\Sigma^{n}$ is a $(b,\varepsilon)$-wise independent distribution, if for every distinct $i_{1},\ldots,i_{b}\in[n]$ it holds that $(X_{i_{1}},\ldots,X_{i_{b}})$ is $\varepsilon$-close to the uniform distribution over $\Sigma^{b}$. Given our efficiently generated small biased spaces, we can efficiently generate almost $b$-wise independent sample spaces as well.
@@
For any positive integers $n$, $m\leq n$, and $b\leq m$, and any $\varepsilon>0$, there exists an explicit $(b,\varepsilon)$-wise independent generator $\mathsf{BI}_{b,\varepsilon}\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}\rightarrow[n]^{m}$ with $d=O(b\log n+\log(1/\varepsilon))$. That is, the distribution formed by picking $z\sim U_{d}$ and outputting $\mathsf{BI}_{b,\varepsilon}(z)$ is $(b,\varepsilon)$-wise independent over $[n]^{m}$. Moreover,
@@
    Given $z\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}$, $\mathsf{BI}_{b,\varepsilon}(z)$ is computable in time $\widetilde{O}(n)$.
@@
    Assume that $\theta\in(0,1/4)$ is such that $\varepsilon\leq\theta\cdot n^{-b/2}$. Then, with probability at least $1-2^{-\Omega(\theta b)}$ over $z\sim U_{d}$, $\mathsf{BI}_{b,\varepsilon}(z)$ has at least $m-(1+4\theta)\frac{m^{2}}{n}$ distinct elements.
@@
Let $q=2^{\lceil\log n\rceil}$, and set $\gamma=\varepsilon\cdot 2^{-\frac{b\log q}{2}}$ and $n_{\mathsf{b}}=n\log q$. Let $S_{\mathsf{b}}\subseteq\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n_{\mathsf{b}}}$ denote the $\gamma$-biased set that is guaranteed to us by Corollary 3.3 and let $S\subseteq[q]^{n}$ be the set that corresponds to treating each consecutive $\log q$ bits as an element of $[q]$. [^12] Since we won’t care about optimizing the dependence on $n$, we do not pre-encode using a bounded-independence generator (as in, say, [NN90, AGHP92]). By the XOR lemma, $S$ is $(b,\varepsilon)$-biased over $[q]^{n}$. Moreover, $\log|S|=O(\log(1/\varepsilon)+k\log q+\log n)=O(\log(1/\varepsilon)+b\log n)$, and each element of $S$ can be generated in time
$\mathopen{}\mathclose{{}\left(n_{\mathsf{b}}+\log(1/\gamma)}\right)\cdot\widetilde{O}(\log n_{\mathsf{b}})=\widetilde{O}(n)+\mathopen{}\mathclose{{}\left(b+\log(1/\varepsilon)}\right)\cdot\widetilde{O}(\log n)=\widetilde{O}(n),$
where the last equality follows since we can always assume that $\varepsilon\geq m^{-n}$. Notice that ignoring the last $n-m$ symbols of each element of $S$ still preserves the above properties, which indeed gives rise to an efficiently samplable $(b,\varepsilon)$-wise independent sample space over $[q]^{m}$.
Next, we argue that most samples contain mostly distinct elements. Towards this end, let $X_{1},\ldots,X_{m}$ be our $(b,\varepsilon)$-wise independent distribution $\mathsf{BI}_{b,\varepsilon}(U_{d})$, and let $Z_{i}$ denote the indicator random variable that is $1$ if and only if $X_{i}$ is a duplicate element (namely, there exists $j<i$ such that $X_{i}=X_{j}$). We are looking to bound $\sum_{i\in[m]}Z_{i}$ with high probability.
@@
Assume that $t\leq b/2$ and $\varepsilon\leq\theta q^{-t}$ for some $\theta>0$. Then, for any distinct $i_{1},\ldots,i_{t}\in[m]$, it holds that $\Pr[Z_{i_{1}}=\ldots=Z_{i_{t}}=1]\leq(1+\theta)(m/q)^{t}$.
@@
Fix indices $j_{1},\ldots,j_{t}$, where each $j_{\ell}<i_{\ell}$. The probability that $X_{i_{\ell}}=X_{j_{\ell}}$ for all $\ell\in[t]$ is at most $q^{-t}+\varepsilon\leq(1+\theta)q^{-t}$, since this event depends on at most $2t\leq b$ random variables. Union-bounding over all choices of $j$-s incurs a multiplicative factor of $\prod_{\ell\in[t]}(i_{\ell}-1)\leq m^{t},$ so overall, $\Pr[Z_{i_{1}}=\ldots=Z_{i_{t}}=1]\leq(1+\theta)(m/q)^{t}$. ∎
Now, 3.6 is sufficient to give us good tail bounds (see, e.g., [HH15, Section 3]). In particular, denoting $\mu=(1+\theta)\frac{m}{q}$ there exists a universal constant $c>0$ such that
@@
which implies Item 2 when $n=q$. Finally, we need to argue that we can also handle the case where $n$ is not a power of $2$ (and so $q>n$). In this case, we can take our $\gamma$-biased set to be over $n_{\mathsf{b}}=\lceil\varepsilon^{-1}\log n\rceil n$ bits, and each consecutive $\lceil\varepsilon^{-1}\log n\rceil$ bits are mapped to $[n]$ by simply taking the corresponding integer modulo $n$. The correctness can be found, e.g., in [Rao07]. ∎
Towards introducing our sampler, we will need the following tail bound for $(b,\varepsilon)$-wise independent random variables.
Lemma 3.7 ([XZ24]).
Let $X\sim\Sigma^{m}$ be a $(b,\gamma)$-wise independent distribution, and fix some $\varepsilon>0$. Then, $X$ is also a $(\delta,\varepsilon)$ sampling distribution, where
$\delta=\mathopen{}\mathclose{{}\left(\frac{5\sqrt{b}}{\varepsilon\sqrt{m}}}\right)^{b}+\frac{\gamma}{\varepsilon^{b}}.$
While the error in Item 2 above is small, it is not small enough for us to simply combine Lemmas 3.7 and 3.5, and we will need to do a mild error reduction. We do this via random walks on expanders and discarding repeating symbols, as was also done in [Vad04]. This gives us the following bounded-independence based sampler.
@@
For any positive integers $m\leq n$, any $\delta_{\Gamma}\in(0,1)$, and any constant $\eta\in(0,1)$ such that $m\leq\frac{\eta}{8}n$, there exists an explicit $(\delta_{\Gamma},\varepsilon_{\Gamma}=2\eta)$ sampler $\Gamma\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}\rightarrow[n]^{m}$ with $d=O\mathopen{}\mathclose{{}\left(\frac{\log n}{\log m}\cdot\log\frac{1}{\delta_{\Gamma}}}\right)$, that satisfies the following additional properties.
@@
    Every output of $\Gamma$ contains distinct symbols of $[n]$, and,
@@
    Given $y\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}$, $\Gamma(y)$ is computable in time $\widetilde{O}(n+\log^{2}\frac{1}{\delta_{\Gamma}}\cdot\frac{\log n}{\log m})$.
@@
Set $b$ to be the smallest integer such that $b\log\frac{\eta\sqrt{m}}{5\sqrt{b}}\geq\log\frac{8}{\delta_{\Gamma}}$, set $m^{\prime}=(1+\eta)m$, $\theta=\eta/4$, and $\gamma=\min\mathopen{}\mathclose{{}\left\{{\frac{1}{8}\eta^{b}\cdot\delta_{\Gamma},\theta\cdot n^{-b/2}}}\right\}$. Notice that $b=O\mathopen{}\mathclose{{}\left(\frac{\log(1/\delta_{\Gamma})}{\log m}}\right)$ and $\log\frac{1}{\gamma}=O\mathopen{}\mathclose{{}\left(\frac{\log n}{\log m}\cdot\log\frac{1}{\delta_{\Gamma}}}\right)$. Let
$\mathsf{BI}_{b,\gamma}\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d^{\prime}}\rightarrow[n]^{m^{\prime}}$
be the $(b,\gamma)$-wise independent generator guaranteed to us by Lemma 3.5, with $d^{\prime}=O(b\log n+\log(1/\gamma))=O(\log(1/\gamma))$. By Lemma 3.7, $X=\mathsf{BI}_{b,\varepsilon}(U_{d^{\prime}})$ is a $(\delta_{\mathsf{b}},\eta)$ sampling distribution, where
$\delta_{\mathsf{b}}=\mathopen{}\mathclose{{}\left(\frac{5\sqrt{b}}{\eta\sqrt{m^{\prime}}}}\right)^{b}+\frac{\gamma}{\eta^{b}}\leq\frac{\delta_{\Gamma}}{8}+\frac{\delta_{\Gamma}}{8}\leq\frac{\delta_{\Gamma}}{4}.$
Also, we know from Lemma 3.5 that with probability at least $1-2^{-\Omega(\theta b)}\triangleq 1-p$, each sample from $X$ has at least $m^{\prime}-(1+4\theta)\frac{m^{\prime 2}}{n}\geq m$ distinct symbols, using the fact that $\frac{n}{m}\geq\frac{(1+\eta)^{3}}{\eta}$. Conditioned on seeing at least $m$ distinct symbols, $X$ as a sampling distribution, when we remove the non-distinct elements, has confidence $\frac{\delta_{\Gamma}/4}{1-p}\leq\frac{\delta_{\Gamma}}{2}$ and accuracy $2\eta$ (where the second $\eta$ comes from the fact that $\eta m$ symbols were removed).
Next, in order to improve the probability of sampling a good sequence, let $G=(V=\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d^{\prime}},E)$ be the $D$-regular $\lambda$-spectral expander of Corollary 2.18, instantiated with $\lambda=p$, so $D\leq p^{-c}$ for some universal constant $c$. Write $d=d^{\prime}+\ell^{\prime}$ for $\ell^{\prime}=\ell\cdot\log D$, where $\ell=c_{\ell}\cdot\frac{\log(1/\delta_{\Gamma})}{b}$ for some constant $c_{\ell}$ soon to be determined. Given $y=(z,w)\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d^{\prime}}\times[D]^{\ell}$, let $z=v_{0},v_{2},\ldots,v_{\ell}$ denote the corresponding random walk over $G$. Our sampler $\Gamma$, on input $y$, computes $\mathsf{BI}_{b,\gamma}(v_{i})$ and outputs the first sequence with at least $m$ distinct symbols. If no such sequence was found, $\Gamma$ simply outputs $(1,\ldots,m)$ (in which case we say it failed). By the expander hitting property (see, e.g., [Vad12, Section 4]), $\Gamma$ fails with probability at most

### v2

@@ 3.1 Fast Generation of Small-Bias Sets
A distribution $S\sim\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}$ is $\varepsilon$-biased if it is indistinguishable from uniform by every linear test. Namely, if for every nonempty $T\subseteq[n]$ it holds that $\Pr_{s\sim S}[\bigoplus_{i\in T}s_{i}=1]\in[\frac{1-\varepsilon}{2},\frac{1+\varepsilon}{2}]$. We say that a set $S\subseteq\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}$ is $\varepsilon$-biased if the flat distribution over its support is $\varepsilon$-biased. We say that a linear code $\mathcal{C}\subseteq\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}$ is $\varepsilon$-balanced if the Hamming weight of each nonzero codeword lies in $\frac{1-\varepsilon}{2}n,\frac{1+\varepsilon}{2}n]$. It is known that these two objects are essentially the same: $S$ is $\varepsilon$-biased if and only if the $|S|\times n$ matrix whose rows are the elements of $S$ is a generator matrix of an $\varepsilon$-balanced code.
@@
Lemma 3.1 (RWs amplify bias [TA-17, Theorem 3.1] [^14] The argument for $t=2$ was suggested already by Rozenman and Wigderson (see [BOG12]). Note that the goal of the RW is to reduce the bias, $\varepsilon\ll\varepsilon_{0}$, but we choose to use “amplify” and follow the existing nomenclature.).
@@
There exist constants $R\in(0,1)$ and $\varepsilon_{0}\in(0,1)$ such that there exists an explicit family of codes $\mathopen{}\mathclose{{\left\{{\mathrm{Jus}_{n}}}}\right\}$ parameterized by block length $n$, with rate $R$ and $\varepsilon_{0}$-balanced. Moreover, given $x\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{k=Rn}$, $\mathrm{Jus}_{n}(x)$ can be computed in $\widetilde{O}(n)$.
@@
The parameters of the codes follow from the original construction (and specifically, the lemma holds, say, with $R=\frac{1}{8}$ and $\varepsilon_{0}=\frac{37}{40}$), so we will just show that the code is efficiently computable. Given a message $x$, we first encode it with a full Reed–Solomon code of constant relative distance over a field $\mathds{F}_{q}$ of characteristic $2$, where $q\log q=O(n)$. By lemma˜2.1, this can be done in time $\widetilde{O}(q)=\widetilde{O}(n)$. Then, we replace each Reed–Solomon symbol $p_{x}(\alpha)$, for $\alpha\in\mathds{F}_{q}$, with the binary representation of $(p(\alpha),\alpha\cdot p(\alpha))$. (In other words, we concatenate Reed–Solomon with the Wozencraft ensemble.) This takes $\widetilde{O}(q)$ time as well. ∎
@@
There exist a constant $c>1$, and an explicit family of balanced codes, such that for every $\bar{n}\in\mathbb{N}$ and any $\varepsilon>0$, $\mathcal{C}\subseteq\mathds{F}_{2}^{\bar{n}}$ is $\varepsilon$-balanced of rate $R=\varepsilon^{c}$, and given $x\in\mathds{F}_{2}^{k=R\bar{n}}$, any $m$ bits of $\mathcal{C}(x)$ can be computed in time $\widetilde{O}(k)+m\log(1/\varepsilon)\cdot\operatorname{polylog}(k)$.
Moreover, for every $k\in\mathbb{N}$ and any $\varepsilon>0$ there exists an explicit $\varepsilon$-biased set over $\mathds{F}_{2}^{k}$ generated by a function $\mathsf{SmallBias}\colon[\bar{n}]\rightarrow\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{k}$ computable in time $(k+\log(1/\varepsilon))\cdot\operatorname{polylog}(k)$.
@@
Let $\mathcal{C}_{0}\colon\mathds{F}_{2}^{k=\Theta(n)}\rightarrow\mathds{F}_{2}^{n}$ be the $\varepsilon_{0}$-balanced code guaranteed to us by lemma˜3.2, and let $G=(V=[n],E)$ be the $D$-regular $\lambda$-spectral expander of corollary˜2.17, instantiated with $\lambda=\frac{1-\varepsilon_{0}}{4}$ (so $D=D(\varepsilon_{0})$). Letting $\mathcal{C}\colon\mathds{F}_{2}^{k}\rightarrow\mathds{F}_{2}^{\bar{n}}$ be the amplified code of lemma˜3.1 set with
$t=\frac{2\log(1/\varepsilon)}{\log\mathopen{}\mathclose{{\left(\frac{2}{1+\varepsilon_{0}}}}\right)}=O(\log(1/\varepsilon)),$
the lemma tells us that it is $(\varepsilon_{0}+2\lambda)^{t/2}\leq\varepsilon$ balanced. Given $x\in\mathds{F}_{2}^{n}$ and $i\in[\bar{n}]$, computing $\mathcal{C}(x)_{i}$ amounts to XORing $t$ coordinates of $\mathcal{C}_{0}(x)$ determined by $i=(v,i_{1},\ldots,i_{t})$, which indexes a random walk over $G$. Computing $\mathcal{C}_{0}(x)$ takes $\widetilde{O}(n)$ time, and computing a length-$t$ walk over $G$ takes $t\cdot\log(1/\lambda)\cdot\operatorname{polylog}(n)$ time. The corollary then follows, observing that $\bar{n}=n\cdot D^{t}=n\cdot\operatorname{poly}(1/\varepsilon)$.
@@
$t\cdot\log(1/\lambda)\cdot\operatorname{polylog}(n)+k\cdot\widetilde{O}(\log n)=(k+\log(1/\varepsilon))\cdot\operatorname{polylog}(k),$
recalling that $k=\Theta(n)$. ∎
@@
Instead of using Justesen’s code from lemma˜3.2 as our inner code $\mathcal{C}_{0}$, we can instead use the linear-time encodable code of Spielman [SPI96]. While not stated as balanced codes, but rather as constant relative distance codes, one can verify that the distance can also be bounded by above. The construction is more involved than Justesen’s one. However, in the logarithmic-cost RAM model, in which basic register operations over $O(\log n)$ bit registers count as a single time step, Spielman’s code can be implemented in $O(n)$ time.
@@ 3.2 A Sampler from Bounded Independence
Recall that $X_{1},\ldots,X_{n}\sim\Sigma^{n}$ is a $b$-wise independent distribution, if for every distinct $i_{1},\ldots,i_{b}\in[n]$ it holds that $(X_{i_{1}},\ldots,X_{i_{b}})=U_{\Sigma^{b}}$, the uniform distribution over $\Sigma^{b}$.
@@
For any prime power $n$, any $m\leq n$, and any $b\leq m$, there exists an explicit $b$-wise independent generator $\mathsf{BI}_{b}\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}\rightarrow[n]^{m}$ with $d=O(b\log n)$ [^16] That is, the distribution formed by picking $z\sim U_{d}$ and outputting $\mathsf{BI}_{b}(z)$ is $b$-wise independent over $n]^{m}$. that satisfies the following.
@@
    Given $z\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}$, $\mathsf{BI}_{b}(z)$ is computable in time $\widetilde{O}(n)$.
@@
    For any $\theta>0$ the following holds. With probability at least $1-2^{-\Omega(\theta b)}$ over $z\sim U_{d}$, $\mathsf{BI}_{b}(z)$ has at least $m-(1+\theta)\frac{m^{2}}{n}$ distinct elements.
@@
We use the standard polynomial-based construction of bounded independence sample spaces. Concretely, given $z\in\mathds{F}_{n}^{b}$, we interpret it as a polynomial $f_{z}(x)=\sum_{i=1}^{b}z_{i}x^{i-1}$, and we let $\mathsf{BI}_{b}(z)$ output $(f_{z}(\alpha_{1}),\ldots,f_{z}(\alpha_{m}))$, where the $\alpha_{i}$-s are distinct elements in $\mathds{F}_{n}$. This gives us a $b$-wise generator over $n]^{m}$ (the correctness can be found, e.g., in [VAD12, Chapter 3]). By item˜1 of lemma˜2.1, $\mathsf{BI}_{b}(z)$ is computable in time $\widetilde{O}(n)$ (and this is true for any $b\leq n$). The seed length $d$ is given by $O(b\log n)$.
Next, we argue that most samples contain mostly distinct elements. Towards this end, let $X_{1},\ldots,X_{m}$ be our $b$-wise independent distribution $\mathsf{BI}_{b}(U_{d})$, and let $Z_{i}$ denote the indicator random variable that is $1$ if and only if $X_{i}$ is a duplicate element (namely, there exists $j<i$ such that $X_{i}=X_{j}$). We are looking to bound $\sum_{i\in[m]}Z_{i}$ with high probability. This will follow from the fact that $X$ is $b$-wise independent.
@@
Assume that $t\leq b/2$. Then, for any distinct $i_{1},\ldots,i_{t}\in[m]$, it holds that $\Pr[Z_{i_{1}}=\ldots=Z_{i_{t}}=1]\leq(m/n)^{t}$.
@@
Fix indices $j_{1},\ldots,j_{t}$, where each $j_{\ell}<i_{\ell}$. The probability that $X_{i_{\ell}}=X_{j_{\ell}}$ for all $\ell\in[t]$ is at most $n^{-t}$, since this event depends on at most $2t\leq b$ random variables. Union-bounding over all choices of $j$-s incurs a multiplicative factor of $\prod_{\ell\in[t]}(i_{\ell}-1)\leq m^{t},$ so overall, $\Pr[Z_{i_{1}}=\ldots=Z_{i_{t}}=1]\leq(m/n)^{t}$. ∎
Now, ˜3.6 is sufficient to give us good tail bounds (see, e.g., [HH15, Section 3] [^17] In the notation of [HH15], the distribution $Z$ is $(0,b)$-growth-bounded. The tail bound then follows from [HH15, Theorem 3.2].). In particular, denoting $\mu=\frac{m}{n}$ there exists a universal constant $c>0$ such that
@@
which implies item˜2. ∎
Towards introducing our sampler, we will need the following tail bound for $b$-wise independent random variables.
Lemma 3.7 (e.g., [VAD12, Chapter 3]).
Let $X\sim\Sigma^{m}$ be a $b$-wise independent distribution, and fix some $\varepsilon>0$. Then, $X$ is also a $(\delta,\varepsilon)$ sampling distribution, where $\delta=\mathopen{}\mathclose{{\left(\frac{b}{2\varepsilon\sqrt{m}}}}\right)^{b}.$
While the error in item˜2 above is small, it is not small enough for us to simply combine lemmas˜3.7 and 3.5, and we will need to do a mild error reduction. We do this via random walks on expanders and discarding repeating symbols, as was also done in [VAD04, Section 8]. This gives us the following bounded-independence based sampler.
@@
For any positive integers $m\leq n$, any $\delta_{\Gamma}\in(0,1)$, and any constant $\eta\in(0,1)$ such that $m\leq\frac{\eta}{8}n$, there exists an explicit $(\delta_{\Gamma},\varepsilon_{\Gamma}=2\eta)$-averaging sampler $\Gamma\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}\rightarrow[n]^{m}$ with $d=O\mathopen{}\mathclose{{\left(\frac{\log n}{\log m}\cdot\log\frac{1}{\delta_{\Gamma}}}}\right)$, that satisfies the following additional properties.
@@
    Every output of $\Gamma$ contains distinct symbols of $[n]$, and
@@
    Given $y\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}$, $\Gamma(y)$ is computable in time $\widetilde{O}(n)+\operatorname{poly}\mathopen{}\mathclose{{\left(\log\frac{1}{\delta_{\Gamma}}\cdot\frac{\log n}{\log m}}}\right)$.
@@
Set $b$ to be the smallest integer such that $b\log\frac{2\eta\sqrt{m^{\prime}}}{b}\geq\log\frac{4}{\delta_{\Gamma}}$ and set $m^{\prime}=(1+\eta)m$, $\theta=\eta/2$. Notice that $b=O\mathopen{}\mathclose{{\left(\frac{\log(1/\delta_{\Gamma})}{\log m}}}\right)$. Assume first that $n$ is a prime power, and let
$\mathsf{BI}_{b}\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d^{\prime}}\rightarrow[n]^{m^{\prime}}$
be the $b$-wise independent generator guaranteed to us by lemma˜3.5, with $d^{\prime}=O(b\log n)$. By lemma˜3.7, $X=\mathsf{BI}_{b}(U_{d^{\prime}})$ is a $(\delta_{\mathsf{b}},\eta)$ sampling distribution, where
$\delta_{\mathsf{b}}=\mathopen{}\mathclose{{\left(\frac{b}{2\eta\sqrt{m^{\prime}}}}}\right)^{b}\leq\frac{\delta_{\Gamma}}{4}.$
Also, we know from lemma˜3.5 that with probability at least $1-2^{-\Omega(\theta b)}\triangleq 1-p$, each sample from $X$ has at least $m^{\prime}-(1+\theta)\frac{m^{\prime 2}}{n}\geq m$ distinct symbols, using the fact that $\frac{n}{m}\geq\frac{(1+\theta)(1+\eta)^{2}}{\eta}$. Conditioned on seeing at least $m$ distinct symbols, $X$ as a sampling distribution, when we remove the non-distinct elements, has confidence $\frac{\delta_{\Gamma}/4}{1-p}\leq\frac{\delta_{\Gamma}}{2}$ and accuracy $2\eta$ (where the second $\eta$ comes from the fact that $\eta m$ symbols were removed).
Next, in order to improve the probability of sampling a good sequence to match the confidence, let $G=(V=\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d^{\prime}},E)$ be the $D$-regular $\lambda$-spectral expander of corollary˜2.17, instantiated with $\lambda=p$, so $D\leq p^{-c}$ for some universal constant $c$. Write $d=d^{\prime}+\ell^{\prime}$ for $\ell^{\prime}=\ell\cdot\log D$, where $\ell=c_{\ell}\cdot\frac{\log(1/\delta_{\Gamma})}{b}$ for some constant $c_{\ell}$ soon to be determined. Given $y=(z,w)\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d^{\prime}}\times[D]^{\ell}$, let $z=v_{0},v_{1},\ldots,v_{\ell}$ denote the corresponding random walk over $G$. Our sampler $\Gamma$, on input $y$, computes $\mathsf{BI}_{b}(v_{i})$ for every $i\in[\ell]$ and outputs the first $m$ distinct symbols of the first sequence with at least $m$ distinct symbols. If no such sequence was found, $\Gamma$ simply outputs $(1,\ldots,m)$ (in which case we say it failed). By the expander hitting property (see, e.g., [VAD12, Chapter 4]), $\Gamma$ fails with probability at most